# Reinforcing Infrastructure Networks with Multicriteria Portfolio Decision Analysis: An Application to Railway Stations in Finland

### v1

@@ 3.3 Fortifying the Networks
$p(x\mid q)=\prod_{k=1}^{n}\left[x_{k}p_{k}(q)+(1-x_{k})(1-p_{k}(q))\right]\in[0,1].$ (3)
This formulation is general in that the probability of the disruption of each node can depend on the entire portfolio of implemented actions. Still, it assumes that the disruption events are independent in that the probability of disruption at a given node is not impacted by whether or not there has been a disruption at some other node(s). Thus, if two actions affect the probability of disruption of a given node and can be taken jointly, it would be necessary to estimate the disruption probability of the node if both actions are implemented or if only one of the actions is taken.
When the DM seeks to maximize the expected network performance, the objective is to determine which feasible portfolios outperform others at different cost levels for all feasible weights. We use the concept of dominance to compare the portfolios.
@@
A portfolio $q^{1}\in\mathcal{Q}_{F}$ is dominated by portfolio $q^{2}\in\mathcal{Q}_{F}$ in the information set $\mathcal{S}$, denoted by $q^{2}\overset{\mathcal{S}}{\succ}q^{1}$, if and only if $\mathbb{E}\left[u(x,w)\mid q^{1}\right]\leq\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ for all $w\in\mathcal{S}$ and (ii) $\mathbb{E}\left[u(x,w)\mid q^{1}\right]<\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ for some $w\in\mathcal{S}$.
The dominance between two portfolios can be determined by comparing the expected utility resulting from implementing them at the extreme points $w_{e}\in\mathcal{S}^{ext}$ of the information set $\mathcal{S}$ [35]. The criterion is formalized in Definition 2. These extreme points of a polyhedral set can be computed using linear programming techniques [see, e.g., 36].
@@
$q^{2}\overset{\mathcal{S}}{\succ}q^{1}\iff\min_{w_{e}\in\mathcal{S}^{\text{ext}}}\mathbb{E}\left[u(x,w_{e})\mid q^{2}\right]-\mathbb{E}\left[u(x,w_{e})\mid q^{1}\right]\geq 0$
and $\text{and }\max_{w_{e}\in\mathcal{S}^{\text{ext}}}\mathbb{E}\left[u(x,w_{e})\mid q^{2}\right]-\mathbb{E}\left[u(x,w_{e})\mid q^{1}\right]>0.$
Moreover, if $\mathbb{E}\left[u(x,w)\mid q^{1}\right]=\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ for all $w\in\mathcal{S}$, then portfolios $q^{1}$ and $q^{2}$ have equal expected performance, which is denoted by $q^{1}\overset{\mathcal{S}}{\sim}q^{2}$. A feasible portfolio is cost-efficient if (i) it is not dominated by any feasible portfolio of equal or lower cost and (ii) there is no other portfolio with lower cost that offers equal performance.
@@
A portfolio $q^{1}\in Q_{F}$ is cost-efficient with respect to another portfolio $q^{2}\in Q_{F}$ in the information set $\mathcal{S}$, denoted by $q^{1}\overset{\mathcal{S}}{\succ_{C}}q^{2}$, if and only if (i) $q^{1}\overset{\mathcal{S}}{\succ}q^{2},c(q^{1})\leq c(q^{2})$ or (ii) $q^{1}\overset{\mathcal{S}}{\sim}q^{2}$ and $c(q^{1})<c(q^{2})$.
Definition 4.
A portfolio $q^{1}\in Q_{F}$ is cost-efficient in the information set $\mathcal{S}$, denoted by $q^{1}\in Q_{CE}\subseteq Q_{F}$ if and only if $\nexists q^{2}\in Q_{F}$ such that $q^{2}\overset{\mathcal{S}}{\succ_{C}}q^{1}$.
Although Definitions 1, 3, and 4 refer to maximization of expected network performance, these definitions can be extended to account for feasibility constraints that arise from the consideration of risk measures such as VaR and CVaR, for example.
@@ 3.4.1 Algorithm to Determine Cost-efficient Portfolios
The Algorithm 1 is proposed for determining cost-efficient portfolios $\mathcal{Q}_{CE}$. The algorithm generates new portfolios at each iteration by adding a reinforcement action to the portfolios in the set of cost-efficient portfolios and the basic portfolios (i.e., portfolios that are not cost-efficient but can potentially be extended to cost-efficient portfolios by adding reinforcement actions.). The set of cost-efficient portfolios is updated at each iteration, and the portfolios that are guaranteed not to be basic are removed, avoiding the complete enumeration of feasible portfolios.
@@
In Step 1, the utility function is computed for all network states and extreme points of the set of feasible weights. If the computation of the utility function is time-consuming, a subset of the states can be evaluated. In Step 2, the set of cost-efficient portfolios is initialized by including only the portfolio with no reinforcement actions. In Step 3, the set of basic portfolios is initialized as an empty set. In Steps 4 to 12, the index $l=1,\dots,h$ iterates through all reinforcement actions. In Step 5, a set $\mathcal{Q}^{l}$ is constructed by taking every portfolio in the set $\mathcal{Q}^{l-1}\cup\mathcal{Q}_{B}^{l-1}$ and modifying its $l^{th}$ action. In Step 6, the portfolios in $\mathcal{Q}^{l}$ are compared to portfolios in $\mathcal{Q}^{l-1}$, and cost-inefficient portfolios are stored in the set $\mathcal{Q}^{1}_{D}$. In Step 7, the cost-inefficient portfolios in $\mathcal{Q}^{1}_{D}$ are removed from the set $\mathcal{Q}^{l}$. In Step 8, the portfolios in $\mathcal{Q}^{l-1}$ are compared to portfolios in $\mathcal{Q}^{l}$, and cost-inefficient portfolios are stored in the set $\mathcal{Q}^{2}_{D}$. In Step 9, the cost-inefficient portfolios in $\mathcal{Q}^{2}_{D}$ are removed from the set $\mathcal{Q}^{l-1}$. In Step 10, the set $Q^{l}$ is updated, including the cost-efficient portfolios from $Q^{l-1}$. In Step 11, the set of basic portfolios is updated, removing the ones that cannot be extended to cost-efficient portfolios. This is done by extending the basic and cost-inefficient portfolios identified in the iteration by including all the remaining reinforcement actions. If the extended portfolio $q^{a}$ of a portfolio $q$ is not cost-efficient, then $q$ is not in the set of basic portfolios. The extended portfolios $q^{a}$ provide an upper bound of the expected utility that can be achieved by extending a portfolio. Still, they are not necessarily feasible (e.g., their cost can be higher than the budget). If a portfolio $q^{a}$ is infeasible, the provided upper bound does not accurately reflect the true value of the maximum expected utility. As a result, the algorithm can keep portfolios in the set basic portfolios that should be discarded. In the A, we provide one complementary approach that can generate tighter bounds, improving the removal of portfolios. The algorithm terminates in Step 13, returning the set of cost-efficient portfolios $Q^{h}$.
3.4.2 Binary Utility Function for Single Objectives
@@
Binary utility functions such as (4) can represent the performance of networks regarding objectives associated with fulfilling a certain condition. For example, in a transportation network, the DM may be interested in having a certain number of edges connected to a specific node for evacuation purposes or in having a connection between two nodes to transport sensitive cargo, such that the objective is either met or not.
@@
$u_{j}(x)=\begin{cases}1,\;\text{if the objective $j$ is met in the network state $x\in\mathcal{X}$.}\\ 0,\;\text{if not.}\end{cases}$ (4)
@@
The expected value of the binary utility function (4) associated with the objective $j$ is equivalent to the probability of achieving the objective.
@@
3.4.3 Minimum Performance Requirements for Single Objectives
The probability of achieving objective $j$ can be enforced to be at least $\alpha_{j}$ by the constraint $\mathbb{E}[u_{j}(x)]\geq\alpha_{j}$. We denote by $\alpha=(\alpha_{1},\ldots,\alpha_{m})$ the vector of all minimum performance requirements, where $\alpha_{j}=0$ if there is no requirement for the single objective $j$.
In general, the Algorithm 1 for determining cost-efficient portfolios does not consider minimum performance requirements for single objectives. This is because it does not consider the utility functions representing the single objectives directly, but the utility function of the network performance (1), which is a weighted sum of them based on the preferences of the DM.
The non-informative set $S^{0}$ is a special case, defined in (2). In this case, there are no stated preferences regarding the different objectives. The extreme points of the set $S^{0}$ are the canonical vectors $e_{j}=(w_{1},\ldots,w_{j},\ldots,w_{m})\in S^{0,\,ext}$, where $w_{j}=1$ and $w_{i}=0$ for all $i\neq j$. At every extreme point $e_{j}\in S^{0,\,ext}$, the utility function of the network is equal to the utility function of the objective $j$, i.e., $u(x,e_{j})=u_{j}(x)$. Therefore, for each objective, there is a cost-efficient portfolio in the non-informative set that, when implemented, results in the maximum achievable performance regarding the objective. This is formalized in Lemma 1.
@@ 3.4.3 Minimum Performance Requirements for Single Objectives
$\forall j=1,\ldots,h\;\exists q^{1}\in\mathcal{Q}_{CE}$ in the information set $\mathcal{S}^{0}$ such that $\mathbb{E}[u_{j}(x)\mid q^{1}]\geq\mathbb{E}[u_{j}(x)\mid q^{2}]\;\forall q^{2}\in\mathcal{Q}_{F}\setminus\mathcal{Q}_{CE}$ and $c(q^{2})\leq c(q^{1})$.
@@
We prove Lemma 1 by contradiction. Consider an objective $j$. Let $q^{1}\in\mathcal{Q}_{CE}$ denote the portfolio with the maximum expected performance in $j$. Assume, for the sake of contradiction, that there exists $q^{2}\in\mathcal{Q}_{F}\setminus\mathcal{Q}_{CE}$ such that $\mathbb{E}[u_{j}(x)\mid q^{2}]>\mathbb{E}[u_{j}(x)\mid q^{1}]$. Since $q^{1}$ is chosen as the portfolio with the maximum expected performance in $j$, it follows that $\nexists q\in\mathcal{Q}_{CE}$ such that $\mathbb{E}[u(x,e_{j})\mid q^{2}]<\mathbb{E}[u(x,e_{j})\mid q]$, where $e_{j}\in\mathcal{S}^{0}$ represents a canonical vector.
@@
Moreover, the cost-efficient portfolios in the information set $S$ capture the maximum possible performance concerning the objective $j$ if $e_{j}\in S$. Therefore, the Algorithm 1 for determining the cost-efficient portfolios can be used when there are minimum performance requirements for single objectives, as long as, for every requirement the corresponding extreme point is in the information set ($\alpha_{j}\neq 0\Rightarrow e_{j}\in S^{ext},\;\;\forall j=1,\ldots,m$). After determining the cost-efficient portfolios, the set must be filtered to remove the ones not fulfilling the requirements.
@@ 3.4.4 Algorithm to Determine Cost-efficient Portfolios with Minimum Performance Requirements
The Algorithm 2 is proposed to identify the cost-efficient portfolios that fulfill minimum performance requirements. The idea is to extend the information set to include the extreme points associated with the minimum performance requirements.
@@
In Step 1, the weight set $\mathcal{S}^{\alpha}$ is constructed to contain one canonical vector $e_{j}$ for every minimum performance requirement $\alpha_{j}$. In Step 2, the extended set $\mathcal{S}^{*}$ is obtained by combining $\mathcal{S}^{\alpha}$ with the extreme points of the set representing the DM preferences $\mathcal{S}^{ext}$. In Step 3, the cost-efficient portfolios in the extended information set $\mathcal{S}^{*}$ are determined using the Algorithm 1. In Step 4, the portfolios that do not fulfill the minimum performance requirements are removed. In Step 5, the cost-efficient portfolios in the information set $\mathcal{S}^{ext}$. The set of portfolios $\mathcal{Q}_{CE}^{\alpha}$ is usually considerably smaller than the set of feasible portfolios, so it is possible to use algorithms such as the one proposed by [37] to identify the set $\mathcal{Q}_{CE}$.
@@ 4 Case Study
We illustrate our methodological approaches by considering the Siilinjärvi train station in Northern Savonia, Finland, depicted in Figure 2(a). The schematic representation of the structure of this station is shown in Figure 2(b). Here, the small squares represent the 22 rail switches, which are mechanical devices that allow trains to change from one track to another. Specifically, these switches correspond to the nodes of the network, while the rail segments between the switches correspond to the edges of the network. Nodes A, B, and C are borderline nodes that define the geographical boundaries of the station.
The station serves the rail commuting needs of the municipality of Siilinjärvi as well as a significant amount of passenger and freight traffic from the south, east, and west. A precondition for this through traffic is that the trains can pass through the station. To do that, trains change tracks to go to different destinations. For example, a train from the south can enter the station at borderline node C and change direction to exit the station at borderline node B to go to the east. For the train to be able to perform that maneuver, some of the switches at the station must operate. The switches required to change direction depend on the origin and end direction, so some combinations are still possible, even when some switches fail.
The rail segments connecting switches are much less prone to possible disruptions than switches, which are relatively complex systems composed of numerous components whose functioning may be thwarted due to technical failures or adverse weather conditions. From this perspective, it is of interest to determine which switches are particularly important to ensure that the transportation objectives related to the provision of connections between the adjacent directions can be met.
Some guidance for estimating the probabilities of switch failures can be inferred from the information offered by condition-based monitoring systems [38]. However, in this example, we illustrate the portfolio optimization approach by assuming that switches are equally prone to failure, i.e., the probability of each switch failure is $p_{k}=0.01$, for the imminent period of operational maintenance. Furthermore, the reinforcement actions that can be applied to the switches are of equal cost and also equally effective in that they reduce this probability by 50% to $p^{\prime}_{k}=0.005$. Both assumptions can be readily altered to reflect the prevailing realities better. Yet, the assumption of equal parameters allows for the deriving of results to better understand the relationship between the switches’ relative importance and their relative position within the station.
@@
4.1 Network Performance
We evaluate the network’s performance based on the reliability of the connections between pairs of borderline nodes A, B, and C. As the connections are bidirectional, we name them (A, B), (B, C), and (A, C). A connection (X,Y) is operational if at least one path connects X with Y without disrupted switches. The reliability of a connection is its probability of being operational. For each connection $j$, we consider the binary utility function (4), where the objective is ”connection $j$ is operational.” Then, we combine the utility functions associated with each connection in the additive utility function (1) to measure the network performance.
The task of evaluating the utility function (4) when the objective is to have an operational connection between a pair of borderline nodes is known in the literature as the terminal-pair reliability problem [39]. This problem is NP-hard and cannot, therefore, be solved in polynomial time [40]. It can be solved for small networks by checking if a path exists between the pair of borderline nodes for each network state. Nevertheless, when considering binary states of the $m$ nodes, the number of states of the network is $2^{m}$, which can be large. To reduce complexity, we use a methodology based on the minimum cut upper bound approximation [41].
Nevertheless, there are algorithms for computing the exact reliability efficiently, even for large networks (see [42] for a survey of recent exact algorithms). If the network size cannot be handled by exact algorithms, or the computation time is critical, one could employ other approaches such as decomposition approaches [43], Monte Carlo simulation in combination with neural networks [44] or Dijkstra’s algorithm [45], or deep neural networks [46], among others.

### v2

@@ 3.3 Fortifying the Network
$p(x\mid q)=\prod_{k=1}^{n}\left[x_{k}(1-p_{k}(q))+(1-x_{k})p_{k}(q)\right]\in[0,1].$ (3)
This formulation is general in that the probability of the disruption of each node can depend on the entire portfolio of implemented actions. Thus, if two actions affect the probability of disruption of a given node, and these two actions can be taken jointly, it would be necessary to estimate the disruption probability of this node for the situation in which both actions are implemented or if only one is taken. Still, the formulation considers that the disruption events are independent in that the probability of disruption at a given node is not impacted by whether or not there has been a disruption at some other node(s).
3.4 Non-Dominated and Cost-Efficient Portfolios
When the DM seeks to maximize the expected network performance, the objective is to determine which feasible portfolios outperform others for all feasible weights that represent the importance of the different objectives. For further insights, such analyses can be produced by comparing portfolios based on the concept of dominance at different cost levels that are feasible with the available budget.
@@ 3.4 Non-Dominated and Cost-Efficient Portfolios
Portfolio $q^{1}\in\mathcal{Q}_{F}$ is dominated by portfolio $q^{2}\in\mathcal{Q}_{F}$ in the information set $\mathcal{S}$, denoted by $q^{2}\overset{\mathcal{S}}{\succ}q^{1}$, if and only if $\mathbb{E}\left[u(x,w)\mid q^{1}\right]\leq\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ for all $w\in\mathcal{S}$ and (ii) $\mathbb{E}\left[u(x,w)\mid q^{1}\right]<\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ for some $w\in\mathcal{S}$.
The dominance between two portfolios can be determined by comparing the expected utility of these portfolios at the extreme points $w_{e}\in\mathcal{S}^{ext}$ of the information set $\mathcal{S}$ [36]. Because this information set is a polyhedral set, these points can be computed with linear programming techniques [see, e.g., 37]. Moreover, if $\mathbb{E}\left[u(x,w)\mid q^{1}\right]=\mathbb{E}\left[u(x,w)\mid q^{2}\right]$ $\forall w\in\mathcal{S}$, then portfolios $q^{1}$ and $q^{2}$ have the same expected performance, which is denoted by $q^{1}\overset{\mathcal{S}}{\sim}q^{2}$. A feasible portfolio is cost-efficient if (i) it is not dominated by any feasible portfolio of equal or lower cost and (ii) there is no other portfolio with lower cost with equal expected performance.
@@
Portfolio $q^{1}\in Q_{F}$ is cost-efficient with respect to another portfolio $q^{2}\in Q_{F}$ in the information set $\mathcal{S}$, denoted by $q^{1}\overset{\mathcal{S}}{\succ_{C}}q^{2}$, if and only if (i) $q^{1}\overset{\mathcal{S}}{\succ}q^{2},c(q^{1})\leq c(q^{2})$ or (ii) $q^{1}\overset{\mathcal{S}}{\sim}q^{2}$ and $c(q^{1})<c(q^{2})$.
@@
Portfolio $q^{1}\in Q_{F}$ is cost-efficient in the information set $\mathcal{S}$, denoted by $q^{1}\in Q_{CE}\subseteq Q_{F}$ if and only if $\nexists q^{2}\in Q_{F}$ such that $q^{2}\overset{\mathcal{S}}{\succ_{C}}q^{1}$.
Although Definitions 1, 2, and 3 refer to the maximization of expected network performance measured by the utility function, these definitions can be extended to represent feasibility constraints that may arise from the consideration of risk measures such as value at risk (VaR) and conditional value at risk (CVaR) [38].
@@ 3.5.1 Algorithm to Determine Cost-efficient Portfolios
Algorithm 1 is presented to determine cost-effective portfolios $\mathcal{Q}_{CE}$. At each iteration, it generates new portfolios by adding a fortification action to the previously computed set of cost-efficient portfolios and the basic portfolios (i.e., portfolios that are not cost-efficient but can potentially be extended to cost-efficient portfolios by adding fortification actions). In particular, the portfolios that cannot belong to the set of basic portfolios are removed, which avoids the complete enumeration of feasible portfolios.
@@
In Step 1, the utility function is computed for all network states and extreme points of the set of feasible weights. If the computation of the utility function is time-consuming, a subset of the states can be evaluated. In Step 2, the set of cost-efficient portfolios is initialized by including only the portfolio with no fortification actions. In Step 3, the set of basic portfolios is initialized as an empty set. In Steps 4 to 12, the index $l=1,\dots,h$ iterates through all fortification actions. In Step 5, a set $\mathcal{Q}^{l}$ is constructed by taking every portfolio in the set $\mathcal{Q}^{l-1}\cup\mathcal{Q}_{B}^{l-1}$ and modifying its $l$-th action. In Step 6, the portfolios in $\mathcal{Q}^{l}$ are compared to portfolios in $\mathcal{Q}^{l-1}$, and cost-inefficient portfolios are stored in the set $\mathcal{Q}^{1}_{D}$. In Step 7, the cost-inefficient portfolios in $\mathcal{Q}^{1}_{D}$ are removed from the set $\mathcal{Q}^{l}$. In Step 8, the portfolios in $\mathcal{Q}^{l-1}$ are compared to portfolios in $\mathcal{Q}^{l}$, and cost-inefficient portfolios are stored in the set $\mathcal{Q}^{2}_{D}$. In Step 9, the cost-inefficient portfolios in $\mathcal{Q}^{2}_{D}$ are removed from the set $\mathcal{Q}^{l-1}$. In Step 10, the set $Q^{l}$ is updated, including the cost-efficient portfolios from $Q^{l-1}$. In Step 11, the set of basic portfolios is updated by including the previous basic portfolios and the cost-inefficient portfolios identified in the iteration that can be extended to cost-efficient portfolios. To determine if a portfolio $q$ is included in basic portfolios, the algorithm constructs its extended portfolio $q^{a}$ by adding all the remaining fortification actions to $q$. If the extended portfolio $q^{a}$ of a portfolio $q$ is not cost-efficient, then $q$ is not in the set of basic portfolios. The extended portfolios $q^{a}$ provide an upper bound of the expected utility that can be achieved by extending a portfolio. Still, they are not necessarily feasible (e.g., their cost can be higher than the budget). If a portfolio $q^{a}$ is infeasible, the provided upper bound does not accurately reflect the value of the maximum expected utility. As a result, the algorithm can keep portfolios in the set basic portfolios that should be discarded. In A, we provide one complementary approach that can generate tighter bounds, improving the removal of portfolios. The algorithm terminates in Step 13, returning the set of cost-efficient portfolios $Q^{h}$.
3.5.2 Binary Utility Function for Individual Objectives
@@
Utility functions such as (4) can be used to represent whether or not the performance of the network satisfies a given condition. For example, in a transportation network, the DM may be interested in having more than two operational edges connected to a specific node for evacuation purposes or in having a connection between two nodes to transport sensitive cargo. In such cases, the resulting objective can be represented by the binary utility function
@@
$u_{j}(x)=\begin{cases}1,\;\text{if objective $j$ is met in network state $x\in\mathcal{X}$.}\\ 0,\;\text{otherwise.}\end{cases}$ (4)
@@
The expected value of the binary utility function (4) associated with objective $j$ is the same as the probability of achieving the objective, i.e.,
@@
3.5.3 Minimum Performance Requirements for Individual Objectives
Depending on the context, there may be a requirement such that the expected utility related to binary objective $j$ is greater than or equal to $\alpha_{j}$. This requirement can be represented by the chance constraint $\mathbb{E}[u_{j}(x)]\geq\alpha_{j}$. The vector of such binary performance requirements is denoted by $\alpha=(\alpha_{1},\ldots,\alpha_{m})$, where $\alpha_{j}=0$ if there is no such requirement.
Algorithm 1 for determining cost-efficient portfolios does not consider such minimum performance requirements for individual objectives because it does not consider the utility functions for individual objectives directly. Instead, it uses the overall utility function of the network performance (1), which is a weighted sum of utility functions for the objectives.
The non-informative set $S^{0}$ is a special case, defined in (2). In this case, there are no stated preferences for individual objectives. The extreme points of the set $S^{0}$ are the canonical vectors $e_{j}=(w_{1},\ldots,w_{j},\ldots,w_{m})\in S^{0,\,ext}$, where $w_{j}=1$ and $w_{i}=0$ for all $i\neq j$. At every extreme point $e_{j}\in S^{0,\,ext}$, the utility function of the network is equal to the utility function of the corresponding objective $j$, i.e., $u(x,e_{j})=u_{j}(x)$. Thus, for each objective, there is a cost-efficient portfolio in the non-informative set that gives the maximum performance on this objective. This is formalized in Lemma 1.
@@ 3.5.3 Minimum Performance Requirements for Individual Objectives
$\forall j=1,\ldots,h\;\exists q^{1}\in\mathcal{Q}_{CE}$ in the information set $\mathcal{S}^{0}$ such that $\mathbb{E}[u_{j}(x)\mid q^{1}]\geq\mathbb{E}[u_{j}(x)\mid q^{2}]\;\forall q^{2}\in\mathcal{Q}_{F}\setminus\mathcal{Q}_{CE}$.
@@
We prove Lemma 1 by contradiction. Consider the objective $j$. Let $q^{1}\in\mathcal{Q}_{CE}$ denote the portfolio with the maximum expected performance in $j$. Assume, for the sake of contradiction, that there exists $q^{2}\in\mathcal{Q}_{F}\setminus\mathcal{Q}_{CE}$ such that $\mathbb{E}[u_{j}(x)\mid q^{2}]>\mathbb{E}[u_{j}(x)\mid q^{1}]$. Since $q^{1}$ is chosen as the portfolio with the maximum expected performance in $j$, it follows that $\nexists q\in\mathcal{Q}_{CE}$ such that $\mathbb{E}[u(x,e_{j})\mid q^{2}]<\mathbb{E}[u(x,e_{j})\mid q]$, where $e_{j}\in\mathcal{S}^{0}$ represents a canonical vector.
@@
Moreover, the maximum possible performance on objective $j$ is attained by some cost-efficient portfolio in the information set $S$ that contains the vector $e_{j}\in S$. Therefore, Algorithm 1 for determining the cost-efficient portfolios can be used in the presence of minimum performance requirements for individual objectives, provided that the corresponding extreme point is in the information set ($\alpha_{j}\neq 0\Rightarrow e_{j}\in S^{ext},\;\;\forall j=1,\ldots,m$). After determining the cost-efficient portfolios, the set must be filtered to remove those that do not fulfill requirements on expected performance.
@@ 3.5.4 Algorithm to Determine Cost-efficient Portfolios with Minimum Performance Requirements
Algorithm 2 can be deployed to identify cost-efficient portfolios that fulfill minimum performance requirements. Toward this end, the information set is extended to include all extreme points associated with the minimum performance requirements.
@@
In Step 1, the weight set $\mathcal{S}^{\alpha}$ is constructed to contain one canonical vector $e_{j}$ for every minimum performance requirement $\alpha_{j}$. In Step 2, the extended set $\mathcal{S}^{*}$ is obtained by combining $\mathcal{S}^{\alpha}$ with the extreme points of the set representing the DM’s preferences $\mathcal{S}^{ext}$. In Step 3, the cost-efficient portfolios in the extended information set $\mathcal{S}^{*}$ are determined with Algorithm 1. In Step 4, portfolios that do not fulfill the minimum performance requirements are removed. In Step 5, the cost-efficient portfolios in the information set $\mathcal{S}^{ext}$ are computed. The set of portfolios $\mathcal{Q}_{CE}^{\alpha}$ is usually much smaller than the set of feasible portfolios, so it is possible to use algorithms such as the one proposed by [39] to identify the set $\mathcal{Q}_{CE}$.
@@ 4 Case Study
We illustrate our methodological contributions by considering the Siilinjärvi train station in Northern Savonia, Finland, depicted in Figure 2(a). The network representation of the structure of this station is in Figure 2(b). Here, the small squares represent the 22 rail switches, which are mechanical devices that allow trains to change from one track to another. Specifically, these switches correspond to the nodes of the network, while the rail segments between the switches correspond to the edges of the network. Nodes A, B, and C are the border nodes that define the geographic boundaries of the station.
The station serves the rail commuting needs of the municipality of Siilinjärvi and a significant amount of passenger and freight traffic from the south, east, and west. A precondition for enabling this traffic is that the trains can pass through the station. To do that, trains change tracks to go in different directions. For example, a train from the south can enter the station at border node C and change direction to exit the station at border node B to go to the east. For the train to be able to perform that maneuver, some of the switches at the station must operate. The switches required to change direction depend on the origin and end direction, so some combinations are still possible, even when some switches fail.
The rail segments connecting switches are much less prone to possible disruptions than switches, which are relatively complex systems composed of numerous components whose functioning may be thwarted due to technical failures or adverse weather conditions. From this perspective, it is interesting to determine which switches are particularly important to ensure that the transportation objectives related to providing connections between the three directions defined by the three border nodes can be met.
Some guidance for estimating the probabilities of switch failures can be inferred from the information offered by condition-based monitoring systems [40]. However, in this example, we illustrate the portfolio optimization approach by assuming that switches are equally prone to failure, i.e., the probability of each switch failure is $p_{k}=0.01$, for the imminent period of operational maintenance. Furthermore, the fortification actions that can be applied to switches are equally costly and equally effective in reducing this probability by 50% ($p^{\prime}_{k}=0.005$). Both assumptions can be readily altered to better reflect the prevailing realities. Yet, the assumption of equal parameters allows for deriving results to better understand the relationship between the switches’ relative importance and their relative position within the station.
@@
4.1 Station Performance
We evaluate the performance of the station based on the reliability of the connections between pairs of border nodes A, B, and C. The connections are bidirectional, and we name them (A, B), (B, C), and (A, C). A connection (X,Y) is operational if at least one path connects X with Y without disrupted switches. The reliability of a connection is its probability of being operational. For each connection $j$, there is an associated binary utility function (4) representing the objective ”connection $j$ is operational.” These three utility functions are combined by the additive utility function (1) to measure the overall performance of the station.
The task of evaluating the utility function (4), which represents the objective of having an operational connection between a pair of border nodes, is known in the literature as the terminal-pair reliability problem [41]. This problem is NP-hard and, therefore, cannot be solved in polynomial time [42]. It can be solved for small networks by checking if a path exists between the pair border nodes for each network state. However, when the number of nodes $n$ is large, this is impractical because there are $2^{n}$ network states, which can become a huge number.
To reduce complexity, we use the minimum cut upper bound approximation [43], recognizing that there are algorithms for computing the exact reliability efficiently even for large networks (see [44] for a survey of recent exact algorithms). If the network size cannot be handled by exact algorithms, or the computation time is critical, one can use approaches such as decomposition approaches [45], Monte Carlo simulation in combination with neural networks [46] or Dijkstra’s algorithm [47], or deep neural networks [48], among others.